The director of research and development is testing a new medicine. She wants to know if there is evidence at the 0.1 level that the medicine relieves pain in more than 358 seconds. For a sample of 62 patients, the mean time in which the medicine relieved pain was 366 seconds. Assume the population standard deviation is 21 .

Step 1 of 6: State the null and alternative hypotheses.

Step 2 of 6: Find the value of the test statistic. Round your answer to two decimal places.

Step 3 of 6: Specify if the test is one-tailed or two-tailed.

Step 4 of 6: Find the P-value of the test statistic. Round your answer to four decimal places.

Step 5 of 6: Identify the level of significance for the hypothesis test.

Step 6 of 6: Make the decision to reject or fail to reject the null hypothesis.
Answer #1

classmate Date Page Given population mean lo = 358 seconds Sample size = n - 62 patients Sample mean EX 366 Seconds populaticlassmate Date Page Step 4 P-value = P(Z >= conoce coiloli = P(Z) 3.00) = 0.0014 [ Probability using noilorib biobante-table]